OpenStudy (empty):

I found a fun function who's roots are only all prime numbers if anyone wants to play around with it.

OpenStudy (empty):

\[f(x) = -2+ \sum_{n=1}^\infty \left\lfloor \cos^2 \left( \frac{\pi x}{n} \right) \right\rfloor\] \[f(x)= 0 \iff \text{ x is prime}\]

OpenStudy (empty):

The square is arbitrary, it could have been absolute value sign on the cosine, oh well! https://www.desmos.com/calculator/i8hgrllcfu Really the thing I found is an infinite series for the divisor counting function, and prime numbers have 2 divisors, so that's how they become roots with that -2 heh.

imqwerty (imqwerty):

greatest prime known till date= \(2^{57,885,161}-1\) roots are prime numbers keep scrolling the graph of this function and someday we will get the new greatest prime and become rich

OpenStudy (empty):

Lolol we can do it :P

OpenStudy (empty):

Ok ikram mentioned to me something about a recurrence relation and I kinda found a sorta-kinda one... haha ok so I define this function where we have the upper bound as an argument: \[t\left(x, N\right)=\sum _{n=1}^{N}f\left(\frac{x}{n}\right)^m\] So for a range, we have kinda a recurrence relation thingy: \[\sqrt{N} < x \le N\] \[ t(x,N)=2*t(x,\sqrt{N}) \]

OpenStudy (empty):

Maybe I wrote that wrong, but basically the idea is that after you count all the factors of a number less than the square root of it, there is an equal number above that, so I doubled it. But maybe this doesn't quite work for some reason I don't know haha.